OpenStudy (anonymous):

If f(x) = 3x + 5 and g(x) = x - 1, then g[f(0)] = ?

OpenStudy (anonymous):

\[f(x) = 3x + 5\] \[g(x) = x-1\] \[\text{Now all we have to do is to replace x of g with f(x)}\] \[g(f(x)) = (3x + 5) -1 \] \[\text{when x = 0}\] \[g(f(0)) = 5 - 1 =4\]

OpenStudy (anonymous):

\(\LaTeX \text{ makes math look so beautiful }\)

OpenStudy (anonymous):

\[f(0) = 3*0+5 = 5\] \[g(f(0)) = g(5) = 5-1=4\]

OpenStudy (anonymous):

the answer is 4